OpenStudy (dcolley):

Solve for x. tan 2x- cot 2x=0

OpenStudy (anonymous):

\[ \tan 2x - \cot 2x=0\\ \tan 2x - \frac{1}{\tan 2x}=0\\ \frac{\tan^22x-1}{\tan 2x}=0\\ \tan^22x=1\\ \tan 2x=\pm1\\ \]

OpenStudy (anonymous):

tan2x=cot2x tan2x=1/tan2x tan^2(2x)=1 tan2x=+-1 2x=arctg(+-1) x=1/2arctg(+-1)

OpenStudy (anonymous):

arctan1=0.785398163 arctan(-1) =-0.785398163

OpenStudy (anonymous):

Suppose that 2x= t Then we have : tan t -cot t = 0 Tan t = ( tan t) ^ (-1) tan^2 t = 1 t = k\[pi\] + \[pi\] /4 k=0,1,2,.... 2x = t x = t/2

OpenStudy (kropot72):

Starting with the identity tan 2x = (2 tanx)/(1 - tan^2 x) I obtained a quadratic in terms of x which when solved gives the following solutions for x: x = 67.5 degrees, x = 22.5 degrees. These solutions check out when substituted in the original equation.

OpenStudy (anonymous):

If you just convert to sin and cos u get Sin^2(2x) = Cos^2(2x) We (should) know that sin = cos at pi/4 or 45 degrees etc